OpenStudy (anonymous):

Solve the equation x + 1 / 3 = 2 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

is it \(\large x+\frac{1}{3}=2\) or is it \(\large \frac{x+1}{3}=2\) ???

OpenStudy (anonymous):

the second one